Compare the following alternatives on the basis of Present worth Analysis at an interest rate of 6% per year.|| First Cost, $Petroleum Based Feedstock Inorganic Based Feedstock Year First Cost Annual Cost Salvage Annual Revenues Cash flow, $ Year Firshow the formula  
Answer #1
Petroleum Based Feedstock
Year First Cost Annual cost Salvage Annual Revenue Cash Flow
0 -350000 -350000
1 -130000 300000 170000
2 -130000 300000 170000
3 -130000 300000 170000
4 -130000 300000 170000
5 -130000 300000 170000
6 -350000 -130000 45000 300000 -135000
7 -130000 300000 170000
8 -130000 300000 170000
9 -130000 300000 170000
10 -130000 300000 170000
11 -130000 300000 170000
12 -130000 45000 300000 215000
NPW 882604.13
Inorganic Based Feedstock
Year First Cost Annual cost Salvage Annual Revenue Cash Flow
0 -120000 -120000
1 -60000 290000 230000
2 -60000 290000 230000
3 -60000 290000 230000
4 -120000 -60000 34000 290000 144000
5 -60000 290000 230000
6 -60000 290000 230000
7 -60000 290000 230000
8 -120000 -60000 34000 290000 144000
9 -60000 290000 230000
10 -60000 290000 230000
11 -60000 290000 230000
12 -60000 34000 290000 264000
NPW 1703103.55

As NPW of Inorganic Based Feedstock is higher it should be selected

Showing formula in Excel

Petroleum Based Feedstock
Year First Cost Annual cost Salvage Annual Revenue Cash Flow
0 -350000 =I8+J8+K8+L8
1 -130000 300000 =I9+J9+K9+L9
2 -130000 300000 =I10+J10+K10+L10
3 -130000 300000 =I11+J11+K11+L11
4 -130000 300000 =I12+J12+K12+L12
5 -130000 300000 =I13+J13+K13+L13
6 -350000 -130000 45000 300000 =I14+J14+K14+L14
7 -130000 300000 =I15+J15+K15+L15
8 -130000 300000 =I16+J16+K16+L16
9 -130000 300000 =I17+J17+K17+L17
10 -130000 300000 =I18+J18+K18+L18
11 -130000 300000 =I19+J19+K19+L19
12 -130000 45000 300000 =I20+J20+K20+L20
NPW =NPV(6%,M9:M20)+M8
=M39-M21
Inorganic Based Feedstock
Year First Cost Annual cost Salvage Annual Revenue Cash Flow
0 -120000 =I26+J26+K26+L26
1 -60000 290000 =I27+J27+K27+L27
2 -60000 290000 =I28+J28+K28+L28
3 -60000 290000 =I29+J29+K29+L29
4 -120000 -60000 34000 290000 =I30+J30+K30+L30
5 -60000 290000 =I31+J31+K31+L31
6 -60000 290000 =I32+J32+K32+L32
7 -60000 290000 =I33+J33+K33+L33
8 -120000 -60000 34000 290000 =I34+J34+K34+L34
9 -60000 290000 =I35+J35+K35+L35
10 -60000 290000 =I36+J36+K36+L36
11 -60000 290000 =I37+J37+K37+L37
12 -60000 34000 290000 =I38+J38+K38+L38
NPW =NPV(6%,M27:M38)+M26